Add or substract as indicated. Express your result in simplest form.
2/5w + 10 - 3/2w - 4
Thank You.

Answer by sdmmadam@yahoo.com(530)

2/5w means (2/5)X(1/w) and we will call (1/w) = t for the time being
Then 2/5w =(2/5)t
So, 3/2w = (3/2)X(1/w) = (3/2)t
2/5w + 10 - 3/2w - 4
=(2/5w -3/2w)+(10-4)(grouping like terms)
=(2/5)t-(3/2)t +6
=[2/5-3/2]t+6
=[(2X2-3X5)/(5X2)]t +6 (finding the lcm)
=[(4-15)/10]t +6
=(-11/10)(1/w)+6 (putting t back to our given (1/w))
=6-(11/10w)
Note: After scanning the steps two or three times you will understand that (1/w) can be pulled out as common in two of the terms and the two fractions can be simplified by finding the lcm even orally!